OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0). y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (jackellyn):

find the slope m=y2-y1/x2-x1 y2=0 y1=4 x2=2 x1=-6 then substitute into this eqn y-y1=m(x-x1) m is slope, choose either point for the eqn

OpenStudy (anonymous):

slope is m=(0-4)/(2-(-6)=-4/8=-1/2 so the point slope form is y − 4 = −1/2(x + 6)